(A)= p +ñ 77 cape As² - 136 p=33 n=77-33 =p-c

SO, how many protons, neutrons and electrons are there. I am so lost, Please help me! We are discussing isotopes.
Answer #1

As 3- A mass number z - Atomic number P = number of n = number e = number - charge of of Proton neutron electron Note : - AniFor Age 2- 33 , A-77 * P = 2 1-33 n - A-P (A = Pan) = 77-33 (n = 44] * & - Ptc for gain of electrons ) = 33 +3 charge is adde